OpenStudy (anonymous):

A furniture shop refinishes chairs. Employees use one of two methods to refinish each chair. method one takes 0.5 hours and the material costs $10. method two takes 1.5 hours and the material costs $5. next week they plan to spend 53 hours in labor and $560 dollars in material for refurnishing chairs. How many chairs should they plan to refinish with each method?

OpenStudy (shane_b):

Let x = number of chairs using method 1 Let y = number of chairs using method 2 Then: 53 = 0.5x + 1.5y 560 = 10x + 5y I assume that you know how to solve this system of equations for x and y. You should end up with making 46 chairs using method 1 and 20 chairs using method 2...for a total of 66 chairs.

OpenStudy (anonymous):

okay method uno: Our factors are .5 hours $10 = 53 hours and 560 dollars so .5 ÷ 53= ? and $560 ÷ $10= ? method dos: Our factors are: 1.5 hours $5 = 53 hours and $560 so 1.5 ÷ 53 = ? and $560 ÷ $5 =?

OpenStudy (shane_b):

@rebeccaskell94: based on his previous question, the questions are really asking what number of chairs should be made using each method to make the most number of chairs.

OpenStudy (anonymous):

How many chairs should they plan to refinish with each method?<--- oh yeah that xD oops

OpenStudy (anonymous):

Shane Wins..You Redeemed yourself. Thanks Alot

OpenStudy (shane_b):

No problem...I did the same thing rebecca did the first time...the question should be worded slightly better :)

OpenStudy (anonymous):

Agreed lol

OpenStudy (anonymous):

Blame Algebra lab.. yall have a good night. thanks again